OpenStudy (anonymous):

g(f(x)) and f(f(x)) when f(x)=3x/x-1 g(x)=x/x-3

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

\[g(f(x))=\frac{\frac{3x}{x-1}}{\frac{3x}{x-1}-3}=\frac{3x}{3x-3(x-1)}=\frac{3x}{3}=x\]

OpenStudy (mertsj):

\[f(g(x))=\frac{3\times \frac{x}{x-3}}{\frac{x}{x-3}-1}=\frac{3x}{x-(x-3)}=\frac{3x}{3}=x\]

OpenStudy (mertsj):

So we see that the functions f and g are inverse functions because by definition if f(g(x))=x and g(f(x))=x, then f and g are inverse functions.

OpenStudy (anonymous):

right :) so how about f(f(x))

OpenStudy (mertsj):

Can you try that now?

OpenStudy (anonymous):

3*3x/x-1/3x/x-1

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

??

OpenStudy (mertsj):

You forgot the -1 in the denominator

OpenStudy (mertsj):



OpenStudy (anonymous):